OpenStudy (anonymous):

Find an equation for the ellipse that satisfies the given conditions. Endpoints of minor axis (0, ±6), distance between foci 14

OpenStudy (anonymous):

b=6 2c=14 that is c =7 a^2=b^2+c^2=6^2+7^2=36+49=85 a^2=85 x^2/36+y^2/85=1

OpenStudy (anonymous):

wait...so is it(x^2)/(36+y^2)/(85)?

OpenStudy (anonymous):

nvr mind